OpenStudy (anonymous):

Which of the following equations best represents the problem? You drove for 425 miles at 60 mph. How long did it take you? D = (425)(60) 425 = 60 T 60 = R(425) 60 = 425 T .

OpenStudy (anonymous):

have answer choicers

OpenStudy (anonymous):

v= x/t

OpenStudy (anonymous):

Or in this case, t=d/v

OpenStudy (anonymous):

D = (425)(60) 425 = 60 T 60 = R(425) 60 = 425 T .

OpenStudy (phi):

do you know that distance = speed * time ? I would replace distance with a number they gave you. do the same for the time. which choice matches?

OpenStudy (anonymous):

You went 425 miles at 60 mph. To find the time that took, it's t = d/v That's time = distance/velocity that's time = 425/60